OpenStudy (anonymous):

How would I solve lim x -> 0 sin(3x)/x

OpenStudy (anonymous):

multiply by 1 in the form of 3/3 then you get 3sin(3x)/(3x) and by a change of variable you get 3sin(u)/u and \[\lim_{u \rightarrow 0}\frac{ 3\sin u }{ u }=3\lim_{u \rightarrow 0}\frac{ \sin u }{ u }=3\]

OpenStudy (anonymous):

Wait, can you eleborate on how you got 3 sinu/u?

OpenStudy (anonymous):

let 3x = u then u->0 as x->0.

OpenStudy (anonymous):

u=0, right?

OpenStudy (anonymous):

you want to get your limit in the form sin(whatever)/whatever because as whatever goes to 0, the limit goes to 1

OpenStudy (anonymous):

I know that, but don't you get 3x on the bottom if you multiplied by 3/3?

OpenStudy (amistre64):

\[\sin u=u-\frac{1}{3!}u^3+\frac{1}{5!}u^5\pm...\] \[\sin (3x)=3x-\frac{1}{3!}(3x)^3+\frac{1}{5!}(3x)^5\pm...\] \[\frac1x\sin (3x)=3-\frac{1}{3!}(3^3x^2)+\frac{1}{5!}(3^5x^4)\pm...\] at x=0; it all zeroes out but the constant

OpenStudy (anonymous):

yes you do\[\frac{ 3 }{ 3 }\frac{ \sin 3x }{ x }=\frac{ 3\sin 3x }{ 3x }\]

OpenStudy (anonymous):

What do you do after that?

OpenStudy (anonymous):

I got up to that on my own, then I'm stuck.

OpenStudy (anonymous):

let 3x = u. as x->0, 3x->0 and u->0. thus \[\lim_{x \rightarrow 0}\frac{ 3\sin 3x }{ 3x }=\lim_{u \rightarrow 0}\frac{ 3\sin u }{ u }=3\lim_{u \rightarrow 0}\frac{ \sin u }{ u }\]

OpenStudy (anonymous):

Oh, so you moved both 3's over behind the limit until after all of the simlpification is complete?

OpenStudy (anonymous):

Oh, NVM, I get it.

OpenStudy (anonymous):

You canceled the 3's from each 3x.

OpenStudy (anonymous):

lol, I"m retarded.

OpenStudy (anonymous):

no only the top 3. the bottom 3 stays to make 3x in the denominator... same as the argument of the sine function in the numerator

OpenStudy (anonymous):

Then you canceled the 3's from the each 3x, right?

OpenStudy (anonymous):

Or am I just seeing things?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ 3\sin 3x }{ 3x }=3\lim_{x \rightarrow 0}\frac{ \sin 3x }{ 3x }=3\lim_{u \rightarrow 0}\frac{ \sin u }{ u }\]